# Taming Data Challenges in ML-based Security Tasks: Lessons from Integrating Generative AI

### v1

@@ III-C Related work
Synthetic data for security.   Compared to other domains, limited work in security has used GenAI tools for data augmentation to improve performance. We do not target privacy-preserving data synthesis. Security work using synthetic data can be categorized as follows: (1) Data augmentation using GenAI methods that are tightly integrated with the classifier. This includes work by Jan et al. [49], where they developed a GAN-based data augmentation scheme for web bot detection. Xu et. al [109] developed a VAE-based data augmentation scheme for intrusion detection. VGX [73] use LLMs to generate textual features retrieved from code samples to boost the performance of software vulnerability detection systems. These methods are tightly coupled with the security classifier and are harder to generalize to other tasks. (2) Off-the-shelf use of GenAI models from vision. Instead of using GenAI models designed specifically for tabular data, researchers have applied GAN models from the vision domain for malicious traffic detection [43, 70]. (3) GenAI methods to generate time-series data. Gong et. al [37] designed a GAN to create time series data for website fingerprinting detection. Taylor et. al [95] generate attack data by performing perturbations in the time series sequences for Controller Area Network attacks. STAN[107] is a Gaussian Mixture Model proposed to generate multivariate time series network traffic data for cybersecurity tasks.
(4) Methods like SMOTE. Prior work extensively uses simple linear interpolation strategies to create synthetic data, including the SMOTE family and its variants [18, 58, 45, 59]. Methods like SMOTE are known to perform poorly on high-dimensional tasks [14], common to security. We compare GenAI with SMOTE-based methods.
@@
Synthetic data in other domains.   Computer vision domain has well-established pipelines for data augmentation that are built into popular ML frameworks [1]. This includes image transformation schemes (e.g., cropping, rotation). Qui et al. [79] apply image transformation-based data augmentation to protect DNNs from backdoor attacks in computer vision tasks. Generative models such as GANs have been extensively used to improve performance of image classifiers, in the fields of agriculture [67], medicine [22] and robotics [69]. The networking community has done work to synthetically generate CDN caches [81], BGP network configurations [8] and network traffic [66, 113, 50].
IV Data-challenged Security Tasks
@@
To address Section III-B’s research questions, we select 7 ML-based security classification tasks. This includes 3 binary and 4 multi-class classification tasks. These tasks were selected because they (a) use tabular data for learning, (b) demonstrate degraded classification performance, (c) exhibit one or more of the data challenges identified in Section II, and (d) have publicly released data and source code, allowing us to reproduce their work. All 7 tasks use real-world threat data, and not synthetic data, making our findings applicable to real-world scenarios.
We highlight two more points about these tasks: (1) Most of these are recently studied security tasks, i.e., 6 out of 7 studies were conducted in the last 3 years. (2) We follow the original authors’ recommended classification schemes. Notably, traditional ML models (e.g., tree-based) often outperform deep learning in these security tasks. Recall that improving performance by using more advanced ML schemes is not our focus. Table I summarizes the tasks.
@@
BGP [97].   A binary classification task, in which Autonomous Systems (AS) are classified as either serial hijackers or non-hijackers. Serial hijackers are adversaries that exploit vulnerabilities in the BGP protocol to repeatedly hijack address prefixes over a long period for malicious purposes. Each data sample contains 52 features covering 8 categories, e.g., prefix origination behavior, prefix visibility, longevity of prefix announcements, and address space fragmentation. This is an extremely data-challenged task that suffers from limited or biased training data challenges (a)-(c) (Section II). The dataset includes just 17 serial hijacker samples and 163 benign samples, used to train an Extremely Randomized Trees classifier [36]. The model achieves only a 60.5% F-score on the attack class. We focus our evaluation on this class, since the benign class already yields a high F-score of 98.53%, leaving minimal room for improvement.
Tor [31].   A binary classification task that labels network traffic as either Tor-based malware or benign. This is developed using a 175 sized feature vector based on website fingerprinting features from Haze et al. [44], capturing connection-level traffic characteristics e.g., packet ordering statistics.
@@
The training dataset has 1,183 malware and 6,932 benign samples used to train an XGBoost classifier which achieves only a 43.9% F-score on the attack class. However, like the BGP task, we focus on the attack class since the benign class F-score class is already high 96%. This task suffers from limited or biased training data challenges (a)-(c) (Section II).
IoT [96].    A binary classification task to distinguish between cryptojacking malware and benign applications using features extracted from IoT device network traffic streams.
Each data sample comprises 273 tabular features extracted from time-series using tsfresh [23]. The training set contains 1,116 malicious and 42,750 benign samples. Four classifiers i.e., a Logistic Regression, Gaussian Naive Bayes, Support Vector Machine and K-Nearest Neighbors are trained to report an average performance. We focus on the Logistic Regression model for simplicity, which achieves a macro F-score of 74.16%. This task also suffers from the limited or biased training data challenges (a)-(c) (Section II).
Task  Features Number of samples Class size F-score
@@
TABLE I: Performance and data statistics for the 7 security tasks. F-score is reported after data normalization. * indicates F-score for only malicious class. Test set statistics and macro F-scores across 12 BODMAS months are presented in Table XI (Appendix) and Figure 2, respectively.
@@
Cookie [15].    A multi-class classification task which classifies website cookies into 4 privacy invasion categories (as per GDPR regulations). Each sample comprises 1,689 features that characterize several cookie properties such as name, domain, path, expiration timestamp, and flags such as ’HttpOnly’ and ’HostOnly’. The original dataset contains 89k, 70k, 46k, and 15k samples across the classes, for which the authors emphasize the high human effort in labeling. To reduce computational effort, we use a random sub-sample of 8.4k, 6.6k, 4.4k, and 1.4k samples per class. These samples train an XGBoost classifier, yielding a macro F-score of 71.8%. This task suffers from limited or biased training data challenges (a), (d) (Section II).
MS-Malware [5].    A multi-class classification problem, where malware files are classified into one of 9 classes, originally from a 2015 Kaggle competition [2]. As per the winning team’s approach [5], each data sample comprises 1,804 features representing the outputs of a hex-dump and disassembled files.
@@
Ahmadi et al. [5] achieve a 98% macro F-score with an XGBoost classifier using the entire training set of 10,868 samples, leaving minimal room for improvement. We induce a data-challenged setting by using only ~4% of the dataset (434 samples) for training. An XGBoost classifier trained on the updated dataset achieves a lower macro F-score of 84%. The task suffers from limited or biased training data challenges (a), (b) and (d) (Section II).
@@
BODMAS [110].   A multi-class classification problem where portable executable (PE) files collected over a period of 13 months are classified into one of 40 malware classes. Each sample comprises of 2,831 statistical features extracted using LIEF feature extractor [99]. A Gradient Boosted Decision Tree (GBDT) classifier is trained on the first month’s data and tested on the next 12 months. We observe that only 5 of the 40 classes show significant performance degradation. Thus, our analysis focuses on these 5 classes. The updated training set includes 49 to 367 samples per class, totaling 752 samples. As shown in Figure 2, performance drops notably in months 5 and 6. This task suffers from limited or biased training data challenges (a), (b), and (d), as well as concept drift (Section II).
Figure 2: Macro F-score for BODMAS classifier.
@@
nPrintML [48, 88].    A multi-class classification of network traffic originating from one of 13 operating systems (OS) for intrusion detection. Using the nPrintML framework [48], raw traffic is converted into tabular form, and AutoML [46] suggests a classifier. Each sample includes 4,169 binary features indicating the presence or absence of protocol fields and payloads. The training set contains between 289 and 689 samples per class, totaling 8,271 samples. An AutoML-selected Light Gradient Boosted Machine (LGBM) classifier is trained, achieving a macro F-score of 75.5%. This task faces limited or biased training data challenges (a), (b), and (d) (Section II).
V Data Augmentation Techniques
@@
To perform data augmentation using synthetic data, we use state-of-the-art GenAI tools and propose a new GenAI scheme called Nimai. We also consider traditional (non-GenAI) data augmentation tools.
V-A Augmentation strategy
@@
All the tasks chosen for our analysis have imbalanced class distributions. We use a simple strategy for data augmentation—we balance out the class distribution, i.e., we generate enough synthetic samples for all the minority classes, until they match the size of the largest class. This class-balancing augmentation strategy can be implemented by generating synthetic samples in 2 ways: (1) Class-conditioning: In this case, the data synthesis tool provides an explicit mechanism for conditioning on the class label while generating samples, or we resort to rejection sampling [13], until we obtain enough samples for each class. All the existing GenAI schemes use this method. (2) Sample-conditioning: Sample-conditioned techniques generate synthetic data by sampling in the vicinity of an existing data sample in the training set (i.e., real samples). The generated sample is assigned the same label as the conditioned sample. Our new GenAI scheme, Nimai can perform both sample and class conditioning.
V-B Existing GenAI tools for tabular data
@@
We describe each GenAI scheme used in our study. The implementation details are in Section -C (Appendix).
@@
TVAE [106].   TVAE is based on a Variational Autoencoder and uses a class-conditional sampling technique. Specifically, TVAE comprises an encoder network that maps to a lower-dimensional latent space, and a decoder network that maps back to the feature space. TVAE uses special data processing schemes to handle tabular data challenges. To handle multimodal continuous-valued columns, TVAE uses Variational Gaussian Mixture (VGM) models  [13] to estimate modes and fit a Gaussian mixture. To generate a synthetic sample, a TVAE samples from the latent space and feeds this sample through the decoder. TVAE does not natively support class-conditional sampling, i.e., generating synthetic data from a specific class. As such, we employ rejection sampling [13] to obtain samples for any given class. We disregard another VAE-based generator [27] as it is limited to binary classification.
@@
CTAB-GAN+ [117].   CTAB-GAN+ is based on a Conditional Generative Adversarial Network (cGAN) and uses a class-conditional sampling technique. CTAB-GAN+ belongs to a series of GAN-based techniques proposed for synthetic tabular data generation [106, 40, 116]. Specifically, it comprises generator and discriminator networks as part of a GAN [38]. Akin to image generation, CTAB-GAN+ uses Wasserstein distance with gradient penalty as loss [40] term to improve training stability for underrepresented sample patterns. Also, it uses training-by-sampling method to over-sample rare data points in the training batches for their emphasized learning. To generate a synthetic sample, CTAB-GAN+ randomly samples from a Gaussian distribution, and feeds this sample through the generator. Continuous features are processed using the same approach as used by TVAE, i.e., mode-specific normalization [106]. Unlike TVAE, CTAB-GAN+ supports class-conditional sampling using a class conditional vector to guide the generator. We use CTAB-GAN+ because it outperforms other GAN-based tabular data generators [106, 40, 116].
@@
TabDDPM [57].   TabDDPM leverages a Diffusion model and uses a class-conditional sampling technique. It comprises a network that progressively denoises Gaussian-corrupted samples as part of a reverse Markov process. For generation, it samples randomly from a Gaussian distribution and recursively feeds this sample through the network. TabDDPM performs hyperparameter search for optimal architecture parameters.
@@
TabSyn [115].   TabSyn leverages a VAE in conjunction with a Diffusion model operating in the VAE latent space and uses a class-conditional sampling technique. To generate a synthetic sample, TabSyn draws from a Gaussian distribution and passes it through a Diffusion model to obtain a VAE latent space sample, which is then decoded using the VAE’s decoder. TabSyn does not natively support class-conditional sampling. Like TVAE, we use rejection-sampling to obtain samples for the required class. TabSyn outperforms other Diffusion-based tabular data generators, i.e., CoDi [61] and STaSy [55].
@@
GReaT [16].   GReaT leverages an LLM for tabular data generation. Specifically, GReaT encodes tabular features to natural language text representation for autoregressive language modeling with an LLM. During training, a pretrained LLM (distilGPT-2 [83]) is fine-tuned on the encoded textual data. To generate a synthetic sample, GReaT encodes feature names or values and autoregressively samples the remainder values as text, then decodes it back to the tabular space. In theory, GReaT can generate synthetic samples by conditioning on any field of the feature vector. However, it does not propose a method for targeted conditioning—i.e., generating samples in the vicinity of certain samples. Thus, we condition only on the class field for class-conditioning, and treat GReaT as a class-conditional approach in our evaluation.
@@
REaLTabFormer [92].    REaLTabFormer (like GReaT) fine-tunes a pre-trained LLM (e.g., GPT-2) to generate realistic tabular data. Continuous features are encoded as fixed-length, sub-columnar string tokens mapped to a unique vocabulary. Training employs early stopping based on similarity between a held-out validation set and synthetic samples to prevent overfitting. During generation, output is constrained to generate tokens belonging to the tokens in the training vocabulary, and class-conditional sampling enables controlled data generation. Other LLM-based generators were not considered because CLLM [86] is proposed for datasets under 100 samples, and TabMT [39] did not make the code available.
V-C Traditional data augmentation tools

### v2

@@ 3.3. Related work
Synthetic data for security.   Compared to other domains, limited work in security has used GenAI tools for data augmentation to improve performance. We do not target privacy-preserving data synthesis. Security work using synthetic data can be categorized as follows: (1) Data augmentation using GenAI methods that are tightly integrated with the classifier. This includes work by Jan et al. (jan2020throwing), where they developed a GAN-based data augmentation scheme for web bot detection. Xu et. al (xu2020toward) developed a VAE-based data augmentation scheme for intrusion detection. VGX (nong2024vgx) use LLMs to generate textual features retrieved from code samples to boost the performance of software vulnerability detection systems. These methods are tightly coupled with the security classifier and are harder to generalize to other tasks. (2) Off-the-shelf use of GenAI models from vision. Instead of using GenAI models designed specifically for tabular data, researchers have applied GAN models from the vision domain for malicious traffic detection (hao2021producing; meghdouri2021controllable). (3) GenAI methods to generate time-series data. Gong et. al (gong2022surakav) designed a GAN to create time series data for website fingerprinting detection. Taylor et. al (taylor2016anomaly) generate attack data by performing perturbations in the time series sequences for Controller Area Network attacks. STAN(xu2021stan) is a Gaussian Mixture Model proposed to generate multivariate time series network traffic data for cybersecurity tasks. (4) Methods like SMOTE. Prior work extensively uses simple linear interpolation strategies to create synthetic data, including the SMOTE family and its variants (smote; mc-ccr; adasyn; krawczyk2019radial). Methods like SMOTE are known to perform poorly on high-dimensional tasks (blagus2013smote), common to security. We compare GenAI with SMOTE-based methods.
@@
Synthetic data in other domains.   Computer vision domain has well-established pipelines for data augmentation that are built into popular ML frameworks (pytorch-vision-transforms). This includes image transformation schemes (e.g., cropping, rotation). Qui et al. (qiu2021deepsweep) apply image transformation-based data augmentation to protect DNNs from backdoor attacks in computer vision tasks. Generative models such as GANs have been extensively used to improve performance of image classifiers, in the fields of agriculture (lu2022generative), medicine (chlap2021review) and robotics (ma2022comprehensive). The networking community has done work to synthetically generate CDN caches (sabnis2021tragen), BGP network configurations (bahnasy2020deepbgp) and network traffic (lin2020using; yin2022practical; jiang2024netdiffusion).
4. Data-challenged Security Tasks
@@
To address Section 3.2’s research questions, we select 7 ML-based security classification tasks. This includes 3 binary and 4 multi-class classification tasks. These tasks were selected because they (a) use tabular data for learning, (b) demonstrate degraded classification performance, (c) exhibit one or more of the data challenges identified in Section 2, and (d) have publicly released data and source code, allowing us to reproduce their work. All 7 tasks use real-world threat data, and not synthetic data, making our findings applicable to real-world scenarios. We follow the original authors’ recommended classification schemes. Notably, traditional ML models (e.g., tree-based) often outperform deep learning in these security tasks. Recall that improving performance by using more advanced ML schemes is not our focus. Table 1 summarizes the tasks.
@@
BGP (testart2019profiling).   A binary classification task, in which Autonomous Systems (AS) are classified as either serial hijackers or non-hijackers. Serial hijackers are adversaries that exploit vulnerabilities in the BGP protocol to repeatedly hijack address prefixes over a long period for malicious purposes. Each data sample contains 52 features covering 8 categories, e.g., prefix origination behavior, prefix visibility, longevity of prefix announcements, and address space fragmentation. This is an extremely data-challenged task that suffers from limited or biased training data challenges (a)-(c) (Section 2). The dataset includes just 17 serial hijacker samples and 163 benign samples, used to train an Extremely Randomized Trees classifier (geurts2006extremely). The model achieves only a 60.5% F-score on the attack class. We focus on this class, as the benign class already achieves a 98.53% F-score with little room for improvement.
@@
Tor (dodia2022exposing).   A binary classification task that labels network traffic as either Tor-based malware or benign. This is developed using a 175 sized feature vector based on website fingerprinting features from Haze et al. (hayes2016k), capturing connection-level traffic characteristics e.g., packet ordering statistics. The training dataset has 1,183 malware and 6,932 benign samples used to train an XGBoost classifier which achieves only a 43.9% F-score on the attack class. However, like the BGP task, we focus on the attack class since the benign class F-score class is already high at 96%. This task suffers from limited or biased training data challenges (a)-(c) (Section 2).
Task  Features Number of samples
Class
size
F-score
@@
Table 1. Performance and data statistics for the 7 security tasks. F-score is reported after data normalization. * indicates F-score for only malicious class. Test set statistics and macro F-scores across 12 BODMAS months are presented in Table 13 (Appendix) and Figure 2 respectively.
IoT (tekiner2022lightweight).    A binary classification task to distinguish between cryptojacking malware and benign applications using features extracted from IoT device network traffic streams. Each data sample comprises 273 tabular features extracted from time-series using tsfresh (christ2018time). The training set contains 1,116 malicious and 42,750 benign samples. Four classifiers i.e., a Logistic Regression, Gaussian Naive Bayes, Support Vector Machine and K-Nearest Neighbors are trained to report an average performance. We focus on the Logistic Regression model for simplicity, which achieves a macro F-score of 74.16%. This task also suffers from the limited or biased training data challenges (a)-(c) (Section 2).
@@
Cookie (bollinger2022automating).    A multi-class classification task which classifies website cookies into 4 privacy invasion categories (as per GDPR regulations). Each sample comprises 1,689 features that characterize several cookie properties such as name, domain, path, expiration timestamp, and flags such as ’HttpOnly’ and ’HostOnly’. The original dataset contains 89k, 70k, 46k, and 15k samples across the classes, for which the authors emphasize the high human effort in labeling. To reduce computational effort, we use a random sub-sample of 8.4k, 6.6k, 4.4k, and 1.4k samples per class. An XGBoost classifier is used yielding a macro F-score of 71.9%. This task suffers from limited or biased training data challenges (a), (d) (Section 2).
@@
MS-Malware (ahmadi2016novel).    A multi-class classification problem, where malware files are classified into one of 9 classes, originally from a 2015 Kaggle competition (kaggle_malware_classification). As per the winning team’s approach (ahmadi2016novel), each data sample comprises 1,804 features representing the outputs of a hex-dump and disassembled files. Ahmadi et al. (ahmadi2016novel) achieve a 98% macro F-score with an XGBoost classifier using the entire training set of 10,868 samples, leaving minimal room for improvement. We induce a data-challenged setting by using only ~4% of the dataset (434 samples) for training. An XGBoost classifier trained on the updated dataset achieves a lower macro F-score of 84%. The task suffers from limited or biased training data challenges (a), (b) and (d) (Section 2).
@@
BODMAS (yang2021bodmas).   A multi-class classification problem where port- able executable (PE) files collected over a period of 13 months are classified into one of 40 malware classes. Each sample comprises of 2,831 statistical features extracted using LIEF feature extractor (LIEF). A Gradient Boosted Decision Tree (GBDT) classifier is trained on the first month’s data and tested on the next 12 months. We observe that only 5 of the 40 classes show significant performance degradation. Thus, our analysis focuses on these 5 classes. The updated training set includes 49 to 367 samples per class, totaling 752 samples. As shown in Figure 2, performance drops notably in months 5 and 6. This task suffers from limited or biased training data challenges (a), (b), and (d), as well as concept drift (Section 2).
Figure 2. Macro F-score for BODMAS classifier.:
@@
nPrintML (holland2021new; cicids2017).    A multi-class classification of network traffic originating from one of 13 operating systems (OS) for intrusion detection. Using the nPrintML framework (holland2021new), raw traffic is converted into tabular form, and AutoML (he2021automl) suggests a classifier. Each sample includes 4,169 binary features indicating the presence or absence of protocol fields and payloads. The training set contains between 289 and 689 samples per class, totaling 8,271 samples. An AutoML-selected Light Gradient Boosted Machine (LGBM) classifier is trained, achieving a macro F-score of 75.5%. This task faces limited or biased training data challenges (a), (b), and (d) (Section 2).
5. Data Augmentation Techniques
@@
To perform data augmentation using synthetic data, we use state-of-the-art GenAI tools and propose a new GenAI scheme called Nimai. We also consider traditional (non-GenAI) augmentation tools.
5.1. Augmentation strategy
@@
All the tasks chosen for our analysis have imbalanced class distributions. We use a simple strategy for data augmentation—we balance the class distribution, i.e., we generate enough synthetic samples for all the minority classes, until they match the size of the largest class. This class-balancing augmentation strategy can be implemented by generating synthetic samples in 2 ways: (1) Class-conditioning: In this case, the data synthesis tool provides an explicit mechanism for conditioning on the class label while generating samples, or we resort to rejection sampling (bishop2006pattern), until we obtain enough samples for each class. All existing GenAI schemes use this method. (2) Sample-conditioning: Sample-conditioned techniques generate synthetic data by sampling in the vicinity of an existing sample in the training set (i.e., real samples). The generated sample is assigned the same label as the conditioned sample. Our new GenAI scheme, Nimai can perform both sample- and class-conditioning.
5.2. Existing GenAI tools for tabular data
@@
We describe each GenAI scheme used in our study. The implementation details are in Section C (Appendix).
@@
TVAE (tvae/ctgan).   TVAE is based on a Variational Autoencoder and uses a class-conditional sampling technique. Specifically, TVAE comprises an encoder network that maps samples to a lower dimensional latent space, and a decoder network that maps back to the feature space. TVAE uses special data processing schemes to handle tabular data challenges. To handle multimodal continuous-valued columns, TVAE uses Variational Gaussian Mixture (VGM) models  (bishop2006pattern) to estimate modes and fit a Gaussian mixture. To generate a synthetic sample, a TVAE samples from the latent space and feeds this sample through the decoder. TVAE does not natively support class-conditional sampling, i.e., generating synthetic data from a specific class. As such, we employ rejection sampling (bishop2006pattern) to obtain samples for any given class. We disregard comparison with another VAE-based generator (dai2019generative) as it is limited to binary classification.
@@
CTAB-GAN+ (ctabgan+).   CTAB-GAN+ is based on a Conditional Generative Adversarial Network (cGAN) and uses a class-conditional sampling technique. CTAB-GAN+ belongs to a series of GAN-based techniques proposed for synthetic tabular data generation (tvae/ctgan; gulrajani2017improved; ctabgan). Specifically, it comprises generator and discriminator networks as part of a GAN (goodfellow2014generative). Akin to image generation, CTAB-GAN+ uses Wasserstein distance with gradient penalty as loss (gulrajani2017improved) term to improve training stability for underrepresented sample patterns. Also, it uses training-by-sampling method to over-sample rare data points in the training batches for their emphasized learning. To generate a synthetic sample, CTAB-GAN+ randomly samples from a Gaussian distribution, and feeds this sample through the generator. Continuous features are processed using the same approach as used by TVAE, i.e., mode-specific normalization (tvae/ctgan). Unlike TVAE, CTAB-GAN+ supports class-conditional sampling using a class conditional vector to guide the generator. We use CTAB-GAN+ because it outperforms other GAN-based tabular data generators (tvae/ctgan; gulrajani2017improved; ctabgan).
@@
TabDDPM (tabddpm).   TabDDPM leverages a Diffusion model and uses a class-conditional sampling technique. It comprises a network that progressively denoises Gaussian-noised samples as part of a reverse Markov process. For generation, it samples randomly from a Gaussian distribution and recursively feeds this sample through the network. TabDDPM performs hyperparameter search for optimal architecture parameters.
@@
TabSyn (tabsyn).   TabSyn leverages a VAE in conjunction with a Diffusion model operating in the VAE latent space and uses a class-conditional sampling technique. To generate a synthetic sample, TabSyn draws from a Gaussian distribution and passes it through a Diffusion model to obtain a VAE latent space sample, which is then decoded using the VAE’s decoder. TabSyn does not natively support class-conditional sampling. Like TVAE, we use rejection-sampling to obtain samples for the required class. TabSyn outperforms other Diffusion-based tabular data generators, i.e., CoDi (lee2023codi) and STaSy (kim2022stasy).
@@
GReaT (GreaT).   GReaT leverages an LLM for tabular data generation. Specifically, GReaT encodes tabular features to natural language text representation for autoregressive language modeling with an LLM. During training, a pretrained LLM (distilGPT-2 (sanh2019distilbert)) is fine-tuned on the encoded textual data. To generate a synthetic sample, GReaT encodes feature names or values and autoregressively samples the remainder values as text, then decodes it back to the tabular space. In theory, GReaT can generate synthetic samples by conditioning on any field of the feature vector. However, it does not propose a method for targeted conditioning—i.e., generating samples in the vicinity of certain samples. Thus, we condition only on the class field for class-conditioning, and treat GReaT as a class-conditional approach in our evaluation.
@@
REaLTabFormer (solatorio2023realtabformer).    REaLTabFormer (like GReaT) fine-tunes a pre-trained LLM (e.g., GPT-2) to generate realistic tabular data. Continuous features are encoded as fixed-length, sub-columnar string tokens mapped to a unique vocabulary. Training employs early stopping based on similarity between a held-out validation set and synthetic samples to prevent overfitting. During generation, output is constrained to generate tokens belonging to the tokens in the training vocabulary, and class-conditional sampling enables controlled data generation. Other LLM-based generators were not considered because CLLM (seedatcurated) is proposed for datasets under 100 samples, and TabMT (gulati2024tabmt) did not make their code available.
5.3. Traditional data augmentation tools